a survey found that 10% of americans believe that they have seen a UFO FOR A EXAMPLE OF 10 PEOPLE FIND EACH PROBALILITY THAT 2 OR 3 PEOPLE BELIEVE THAT THEY HAVE SEEN A UFO
Answer #1

here this is binomial distribution with parameter n=10 and p=0.1

PROBALILITY THAT 2 OR 3 PEOPLE BELIEVE THAT THEY HAVE SEEN A UFO

=P(X=2)+P(X=3)=10C2(0.1)2(0.9)8+10C3(0.1)3(0.9)7=0.193710+0.057396=0.251106